OpenStudy (anonymous):

Solve the equation on the interval (0, 2pi): 12 cos^2 x-9=0

myininaya (myininaya):

isolate the cos^2(x) first

OpenStudy (anonymous):

ok so cos^2=9/12 cos^2=3/4?

myininaya (myininaya):

ok you mean cos^2(x)=3/4 cool

myininaya (myininaya):

now take square root of both sides and you will have two equations to solve

OpenStudy (anonymous):

ok ummm hold on

OpenStudy (anonymous):

cosx=sqrt3/2?

myininaya (myininaya):

plus or minus

myininaya (myininaya):

you have \[\cos(x)= \pm \frac{\sqrt{3}}{2}\]

myininaya (myininaya):

\[\cos(x)=\frac{\sqrt{3}}{2} \text{ or } \cos(x)=\frac{-\sqrt{3}}{2}\] we need to solve both of these equations a good place to start is looking at the unit circle

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pi/6?

OpenStudy (anonymous):

and 5pi/6?

myininaya (myininaya):

\[\cos(x)=\frac{\sqrt{3}}{2} \text{ has solutions } x=\frac{\pi}{6} \text{ or also what ?} \\ \cos(x)=\frac{-\sqrt{3}}{2} \text{ has solutions } x=\frac{5\pi}{6} \text{ or also what ? }\]

myininaya (myininaya):

there are two more solutions one for each equation above

OpenStudy (anonymous):

11pi/6 and 7pi/6?

myininaya (myininaya):

\[\cos(x)=\frac{\sqrt{3}}{2} \text{ has solutions } x=\frac{\pi}{6} \text{ or } x=\frac{11 \pi}{6} \\ \cos(x)=\frac{-\sqrt{3}}{2} \text{ has solutions } x=\frac{5\pi}{6} \text{ or } x=\frac{7 \pi}{6}\]

myininaya (myininaya):

yep! :) good job

OpenStudy (anonymous):

yay! so thats my answer?

OpenStudy (anonymous):

but what about the part where is says on the interval 0, 2pi?

myininaya (myininaya):

There are no more solutions in that interval. We already found them all.

OpenStudy (anonymous):

ohhhhh!! thank you! omg thank you so much!!!!